What is the DeBroglie wavelength of an antineutrino with a kinetic energy of 1.8 MeV? Take the mass of the antineutrino to be 4.99E-37 kg.

1.75 nm

2.88E-4 nm

5.36E-16 nm

8.45E-12 nm

1.24 nm
Answer #1

know De Broglie wavelength x= 4/3 = Tome rame Vzx 4.99x1037 6.626 x 10 -34 и x 1.8 x 108x1.6x10 d=1.24x10- 9 x = 1.24 nm (Ans